OpenStudy (anonymous):

can someone help with this? Really really urgent!! Average speed problems!!! this is the first question: What is the value of time, t, for the given values of distance, d, and rate, r? d = 198 km, r = 36 km/h A. 7128 km B. 7128 h C. 5.5 km D. 5.5 h

OpenStudy (anonymous):

i will fan and medal!!!!!

OpenStudy (thatonegirl_):

r=d/t 36=198/t (198)36=t t=7128

OpenStudy (thatonegirl_):

rate is the same thing as speed and speed is always distance over time.

OpenStudy (anonymous):

is it 7128 h or km?

OpenStudy (thatonegirl_):

well we solved for t which is time. You would measure time in hours so 7128h

OpenStudy (anonymous):

ohhhh so you are basically just multiplying ?

OpenStudy (thatonegirl_):

yes I'll write it in an equation, one min.

OpenStudy (anonymous):

can you help with a few more?

OpenStudy (anonymous):

Paul's family drove 377 miles to the beach averaging 58 mi/h on the way there. On the return trip home, they averaged 65 mi/h. What was the total time Paul's family spent driving to and from the beach?

OpenStudy (thatonegirl_):

\[36=\frac{ 198 }{ t }\]

OpenStudy (thatonegirl_):

I just realized I did it wrong. hang on

OpenStudy (anonymous):

A. 11.3 h B. 11.6 h C. 12.3 h D. 13 h

OpenStudy (anonymous):

okay

OpenStudy (thatonegirl_):

\[36t=198\]

OpenStudy (thatonegirl_):

\[t=\frac{ 198 }{ 36 }\]

OpenStudy (anonymous):

ohhhh okay

OpenStudy (thatonegirl_):

so 5.5 hours

OpenStudy (thatonegirl_):

Okay so for the other one... You have a total of 577 miles correct?

OpenStudy (thatonegirl_):

377*